OpenStudy (anonymous):

Solve for 'X' 3-7(x+3)=-4

OpenStudy (anonymous):

3 - 7x - 21 = -4 then do you have any idea?

OpenStudy (anonymous):

Yea is it x=21 or x=-2

OpenStudy (anonymous):

3+4-21 = 7x

OpenStudy (anonymous):

I don't know this mr steve hawking i thought to combine like terms as 3-21-7x=-4 which becomes -18-7x=-4 than +18 added to each sides getting -7x=14 than +7 times both sides or -7 times both sides?

OpenStudy (anonymous):

wait i got this only way to cancle out -7x is to times both sides with -7 leaving x

OpenStudy (anonymous):

wait,you made a mistake!

OpenStudy (anonymous):

-18 - 7x = -4 --> +18 - 18 - 7x = -4 + 18 --> -7x = 14 --> x = -2

OpenStudy (anonymous):

and x = -2 not 21

OpenStudy (anonymous):

ok?